Old Country Links, Inc., produces sausages in three production departments—Mixing, Casing and Curing, and Packaging. In the MRequired 1 Required 2 Required 3 Required 4 Required 5 Determine the Casing and Curing Departments equivalent units of produRequired 1 Required 2 Required 3 Required 4 Required 5 Compute the Casing and Curing Departments cost per equivalent unit foRequired 1 Required 2 Required 3 Required 4 Required 5 Compute the Casing and Curing Departments cost of ending work in procRequired 1 Required 2 Required 3 Required 4 Required 5 Compute the Casing and Curing Departments cost of units transferred oRequired 1 Required 2 Required 3 Required 4 Required 5 Prepare a cost reconciliation report for the Casing and Curing Departm
Answer #1

Solution

1) Equivalent unit

Mixing Material Conversion
Unit transferred out 109 109 109
Ending WIP 7 7*20% = 1.40 7*10% = 0.70
Equivalent unit 116 110.40 109.70

2) Cost per equivalent unit

Mixing Material Conversion
Beginning WIP $13,692 $35 $3,563
Cost added $2,05,548 $19,947 $1,63,181
Total cost $2,19,240 $19,982 $1,66,744
Equivalent unit 116 110.40 109.70
Cost per equivalent unit $1,890 $181 $1,520

3) Cost of ending WIP

Mixing Material Conversion Total
Cost of ending WIP 7*$1890 = $13,230 1.4*$181 = $253 0.7*1$520 = $1,064 $14,547

4) Cost of unit transferred out

Mixing Material Conversion Total
Cost of unit transferred out 109*$1890 = $206,010 109*$181=$19,729 109*$1520=$165,680 $3,91,419

5) Cost reconciliation

Cost accounted for
Beginning WIP $17,290
Cost added $388,676
Total cost accounted for $405,966
Cost accounted as
Cost of unit transferred out $391,419
Cost of ending WIP $14,547
Total cost accounted as $4,05,966

Please give thumbs up if you like my answer :-)